Need help in solving this:
The area in square feet for a rectangular deck is given by A = x^2 + 6x +8

a) What would be the area if x was 6?

b) Remembering that Area = Length * Width, factor the above equation. If the width of
the deck is x + 2, what is the length?

Answer by rapaljer(4671)

a) If x=6, then A=x^2 + 6x +8
A=6^2 +6*6+8
A=36+36+8
A= 80 square units

b) If A=x^2 +6x+8
A=(x+2)(x+4)

A= W*L
If x+2 = width, then L=x+4.

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus